OpenStudy (anonymous):

which of the following are polynomials that are perfect square trinomials: Answer 9x2 - 6x + 1 16x2 + 24x + 9 36b2 - 24b + 8 4a2 - 10a + 25

Directrix (directrix):

A perfect square trinomial is the square of a binomial. To see what happens when we square any binomial, let us square (a + b): (a + b)² = (a + b)(a + b) = a² + 2ab + b² The square of any binomial produces the following three terms: 1. The square of the first term of the binomial: a² 2. Twice the product of the two terms: 2ab 3. The square of the second term: b²

OpenStudy (anonymous):

you want to factor out the coefficient of x^2 from each term and see if half of the "b" term, that is the x term, squared is equal to the third term, or constant term

OpenStudy (anonymous):

can someone just answer it for ,me I want to double check my answers

OpenStudy (anonymous):

which ones do you have?

Directrix (directrix):

9x2 - 6x + 1, 16x2 + 24x + 9

OpenStudy (anonymous):

the last 3

OpenStudy (anonymous):

the first one works because after factoring out a 9 from each term you are left with: 9(x^2-2/3x+1/9) [1/2(2/3)]^2=1/9

OpenStudy (anonymous):

so the first one is right?

OpenStudy (anonymous):

first one is one answer. use the same logic to check the others

OpenStudy (anonymous):

I think second one is right

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and the last one is right?

OpenStudy (anonymous):

yep, 3rd one doesnt work

OpenStudy (anonymous):

actually last one is wrong too, sorry

OpenStudy (anonymous):

wait im confused which one is right 3rd one or last one?

OpenStudy (anonymous):

4(a^2-5/2a+25/4) (5/4)^2 doesn't equal 25/4

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

mmk so the first 2 r right and last 2 r wrong?

OpenStudy (anonymous):

yes

Directrix (directrix):

9x2 - 6x + 1 and 16x2 + 24x + 9